# Synchronization-induced violation of thermodynamic uncertainty relations

### v1

@@ Appendix A System-reservoir interactions
Here we provide some details on the microscopic model of the reservoirs and their couplings with the working medium of the quantum thermal machine. Following the Caldeira-Leggett approach [64], the $\nu$-th bath Hamiltonian reads
@@
with $c_{k,\nu}$ describing the coupling strengths between the QHOs and the $k$-th mode of the $\nu$-th reservoir, modulated by the drives $g_{\nu}^{(l)}(t)$. In the above equation we used the convention according to which if $l=A$ then $\bar{l}=B$, and vice versa. Note that the superscript (t) reminds the time-dependent modulation $g_{1}^{(l)}(t)$. It is important to note that the interaction in Eq. (26) includes counter-term contributions having a twofold purpose: (i) to avoid renormalization of the characteristic frequencies $\omega_{A,B}$ of the QHOs and (ii) to cancel the direct coupling between the latter that would naturally arise [70].
@@ Appendix B Derivation of power $P_{A}$ and associated fluctuations $D_{P_{A}}$
Here we provide details on the derivation of the subsystem average power $P_{l}$ and its fluctuations. We start from the definition in Eq. (5) and we first focus on the term present in the r.h.s.
$\langle P_{l}(t)\rangle={\rm Tr}\bigg{[}\frac{\partial H_{{\rm int},1,l}^{(t)}(t)}{\partial t}\rho(t_{0})\bigg{]}~{},$ (27)
@@
where ${\Gamma}_{1}(t)$ is defined through Eq. (3) via $\gamma_{1}(t)=\frac{1}{m}\theta(t){\Gamma}_{1}(t)$. Here, we have introduced the symmetric and anti-symmetric contributions of the correlation function of Eq. (II.1) $\langle\xi_{\nu}(t)\xi_{\nu^{\prime}}(s)\rangle\equiv{\cal L}_{\nu}(t-s)\delta_{\nu,\nu^{\prime}}$, with ${\cal L}_{\nu}(t-s)={\cal L}_{\nu,{\rm s}}(t-s)+{\cal L}_{\nu,{\rm a}}(t-s)$ where
@@
Recalling Eqs. (29)–(30), we introduce the retarded Green function of the fluctuating force for the $\nu=1$ bath
@@
Using the relations introduced so far, it is possible to prove that real and imaginary part of the Fourier transform of the memory damping kernel $\gamma_{1}(t)$ in Eq. (3) are, respectively,
@@
Now, considering the monochromatic drives of Eq. (2) and following similar steps as done in Ref. [70] we finally arrive at $P_{l}=P_{l}^{(0)}+\delta P_{l}$, where
@@
×$\quad\quad\quad\times\left[{\chi_{2}^{(l,l)}}^{\prime\prime}(\omega)+\cos(\phi){\chi_{2}^{(l,\bar{l})}}^{\prime\prime}(\omega)\right],$ (43)
$\delta P_{A/B}=\mp\Omega\sin(\phi)\int_{-\infty}^{+\infty}\frac{{\rm d}\omega}{4\pi m}\left[\zeta_{1}^{\prime\prime}(\omega+\Omega){\chi_{2}^{(A,B)}}^{\prime}(\omega)\right.$
×$\left.\times\coth\left(\frac{\omega+\Omega}{2T_{1}}\right)\!+\zeta_{1}^{\prime}(\omega+\Omega){\chi_{2}^{(A,B)}}^{\prime\prime}(\omega)\coth\left(\frac{\omega}{2T_{2}}\right)\right].$ (44)
Notice that the total power is eventually given by $P=P_{A}^{(0)}+P_{B}^{(0)}$ and it is reported in the main text in Eq. (15). The expressions for the average heat currents can be derived analogously and are reported in Eq. (II.4). As a final remark we quote the expressions of the fluctuations for the separate power contributions obtained at the lowest perturbative order by following similar steps as outlined above:
$D_{P_{l}}=$ $\Omega^{2}\int_{-\infty}^{+\infty}\frac{{\rm d}\omega}{4\pi m}{\cal J}_{1}(\omega+\Omega)N(\omega,\Omega)$
@@
×$\!\!\times\!\big{[}{\chi_{2}^{(A,A)}}^{\prime\prime}\!(\omega)\!+\!{\chi_{2}^{(B,B)}}^{\prime\prime}\!(\omega)\!+\!2\cos(\phi){\chi_{2}^{(A,B)}}^{\prime\prime}\!(\omega)\big{]}.$ (46)
@@ Appendix C Proof of $Q_{P}\geq 2$
see Eq. (7). We start observing that Eq. (15) can be rewritten as
@@
$\chi_{\mathrm{eff}}^{\prime\prime}(\omega)={\chi_{2}^{(A,A)}}^{\prime\prime}\!(\omega)\!+\!{\chi_{2}^{(B,B)}}^{\prime\prime}\!(\omega)\!+\!2\cos(\phi){\chi_{2}^{(A,B)}}^{\prime\prime}\,,$ (49)
and we can similarly rewrite the expression of the total power fluctuations of Eq. (B). Finally, the entropy production rate $\dot{S}=-J_{1}/T_{1}+(P+J_{2})/T_{2}$ using Eqs. (15)-(II.4) reads
@@
By plugging Eqs. (B), (48) and (50) into Eq. (47) and performing straightforward algebra, the following condition is found:
@@
and where $\lambda_{\omega}=\frac{\omega}{2T_{2}}-\frac{\omega+\Omega}{2T_{1}}$. We will now show that the condition of Eq. (51) can never be satisfied.
@@
Clearly, since $\lambda_{\omega}\coth(\lambda_{\omega})\geq 1$ one has $G(\omega,\omega^{\prime})\geq\left(\lambda_{\omega}-\lambda_{\omega^{\prime}}\right)^{2}\geq 0$ and therefore the sign of Eq. (C) is determined by the other factors in its integrand.
@@
The denominator of Eq. (55) is strictly positive, while the identity
@@
One then concludes that the sign of the integrand in Eq. (C) is given by
@@
$\mathcal{N}(\omega)=$ $2\omega^{4}+2\cos(\phi)(\omega^{2}-\omega_{A}^{2})(\omega^{2}-\omega_{B}^{2})$
-- $2\omega^{2}(\omega_{A}^{2}+\omega_{B}^{2})+\omega_{A}^{4}+\omega_{B}^{4}\,,$ (60)
and where $\mathcal{D}(\omega)$ is given in Eq. (9). Here, $\mathcal{N}(\omega)$ is a quadratic form in $\omega^{2}$ with discriminant $-4(\omega_{A}^{2}-\omega_{B}^{2})^{2}\sin^{2}(\phi)\leq 0$, which proves that $\mathcal{N}(\omega)\geq 0$, whence we conclude that $\omega\chi_{\mathrm{eff}}^{\prime\prime}(\omega)\geq 0$. This finally shows that the integrand of Eq. (C) is non-negative and thus $I\geq 0$, contradicting the condition stated in Eq. (51). Thus, it is proven that $Q_{P}\geq 2$.
@@ Appendix D Asymptotic expressions
$P_{l}\sim\delta P_{l}\to\Omega{\delta\alpha}^{({\rm ad})}_{l},$ (63)
@@
The associated fluctuations start quadratically as
$D^{({\rm ad})}_{P_{l}}$ $=\Omega^{2}\int_{-\infty}^{+\infty}\frac{{\rm d}\omega}{4\pi m}{\cal J}_{1}(\omega)N(\omega,0)$
@@
$P\to\Omega{\cal J}_{1}(\Omega)\sum_{l=A,B}\alpha_{l}^{({\rm dia})},$ (66)
@@
$\delta P_{l}\to\Omega{\cal J}_{1}(\Omega)\left[\delta\alpha_{l}^{({\rm dia})}\frac{\Omega}{\omega_{c}}+\delta\beta^{({\rm dia})}_{l}\right]~{},$ (68)
@@
$\delta\alpha^{({\rm dia})}_{A/B}=\pm\int_{-\infty}^{+\infty}\frac{{\rm d}\omega}{4\pi m}\coth\left(\frac{\omega}{2T_{2}}\right){\chi_{2}^{(A,B)}}^{\prime\prime}(\omega)~{}$
@@
$D^{({\rm dia})}_{P_{l}}\!\!=\!\Omega^{2}{\cal J}_{1}(\Omega)\!\coth\!\!\left(\!\frac{\Omega}{2T_{1}}\!\right)\!\!\int_{-\infty}^{+\infty}\!\!\!\!\frac{{\rm d}\omega}{4\pi m}\!\coth\!\!\left(\!\frac{\omega}{2T_{2}}\!\right)\!{\chi_{2}^{(l,l)}}^{\prime\prime}\!\!(\omega).$ (70)
Author contributions
L. R., M. C., F. C., and M. S. developed the model and performed the calculations. M.S. conceived and together with G. B. supervised the study. All authors discussed the results and contributed to writing and revising the manuscript.
Data availability
Data sets generated during the current study are available from the corresponding author on reasonable request.
Competing interests
The authors declare no competing interests.
Acknowledgements.
M.C. and M.S. acknowledge support from the project PRIN 2022 - 2022PH852L (PE3) TopoFlags - ”Non reciprocal supercurrent and topological transition in hybrid Nb-InSb nanoflags” funded within the programme ”PNRR Missione 4 - Componente 2 - Investimento 1.1 Fondo per il Programma Nazionale di Ricerca e Progetti di Rilevante Interesse Nazionale (PRIN)”, funded by the European Union - Next Generation EU. L.R. and G.B. acknowledge financial support from the Julian Schwinger Foundation (Grant JSF-21-04-0001), from INFN through the project “QUANTUM”, and from the project PRIN 2022 - 2022XK5CPX (PE3) SoS-QuBa - ”Solid State Quantum Batteries: Characterization and Optimization” funded within the programme ”PNRR Missione 4 - Componente 2 - Investimento 1.1 Fondo per il Programma Nazionale di Ricerca e Progetti di Rilevante Interesse Nazionale (PRIN)”, funded by the European Union - Next Generation EU.

### v2

@@ Appendix A System-reservoir interactions
Here we provide some details of the microscopic model of the reservoirs and their couplings with the working medium of the quantum thermal machine. Following the Caldeira-Leggett approach [68], the $\nu$-th bath Hamiltonian reads
@@
with $c_{k,\nu}$ describing the coupling strengths between the QHOs and the $k$-th mode of the $\nu$-th reservoir, modulated by the drives $g_{\nu}^{(l)}(t)$. In the above equation we used the convention according to which if $l=A$ then $\bar{l}=B$, and vice versa. Note that the superscript (t) reminds the time-dependent modulation $g_{1}^{(l)}(t)$. It is important to note that the interaction in equation (28) includes counter-term contributions having a twofold purpose: (i) to avoid renormalization of the characteristic frequencies $\omega_{A,B}$ of the QHOs and (ii) to cancel the direct coupling between the latter that would naturally arise [75].
@@ Appendix B Derivation of power $P_{A}$ and associated fluctuations $D_{P_{A}}$
Here we provide details on the derivation of the subsystem average power $P_{l}$ and its fluctuations. We start from the definition in equation (5) and we first focus on the term present in the r.h.s.
$\langle P_{l}(t)\rangle={\rm Tr}\bigg{[}\frac{\partial H_{{\rm int},1,l}^{(t)}}{\partial t}\rho(t)\bigg{]}~{},$ (29)
@@
where ${\Gamma}_{1}(t)$ is defined through equation (3) via $\gamma_{1}(t)=\frac{1}{m}\theta(t){\Gamma}_{1}(t)$. Here, we have introduced the symmetric and anti-symmetric contributions of the correlation function of equation (II.1) $\langle\xi_{\nu}(t)\xi_{\nu^{\prime}}(s)\rangle\equiv{\cal L}_{\nu}(t-s)\delta_{\nu,\nu^{\prime}}$, with ${\cal L}_{\nu}(t-s)={\cal L}_{\nu,{\rm s}}(t-s)+{\cal L}_{\nu,{\rm a}}(t-s)$ where
@@
Recalling equations (31)–(32), we introduce the retarded Green function of the fluctuating force for the $\nu=1$ bath
@@
Using the relations introduced so far, it is possible to prove that real and imaginary part of the Fourier transform of the memory damping kernel $\gamma_{1}(t)$ in equation (3) are, respectively,
@@
Now, considering the monochromatic drives of equation (2) and following similar steps as done in Ref. [75] we finally arrive at $P_{l}=P_{l}^{(0)}+\delta P_{l}$, where
@@
×$\quad\quad\times\left[f_{l}^{2}(\Omega){\chi_{2}^{(l,l)}}^{\prime\prime}(\omega)+\cos(\phi)f_{l}(\Omega)f_{{\bar{l}}}(\omega){\chi_{2}^{(l,\bar{l})}}^{\prime\prime}(\omega)\right],$ (45)
$\delta P_{A/B}=\mp\Omega f_{A}(\Omega)f_{B}(\Omega)\sin(\phi)\int_{-\infty}^{+\infty}\frac{{\rm d}\omega}{4\pi m}$
×$\quad\quad\times\left[\zeta_{1}^{\prime\prime}(\omega+\Omega){\chi_{2}^{(A,B)}}^{\prime}(\omega)\coth\left(\frac{\omega+\Omega}{2T_{1}}\right)\right.$
$\quad\quad\left.+\zeta_{1}^{\prime}(\omega+\Omega){\chi_{2}^{(A,B)}}^{\prime\prime}(\omega)\coth\left(\frac{\omega}{2T_{2}}\right)\right].$ (46)
Notice that the total power is eventually given by $P=P_{A}^{(0)}+P_{B}^{(0)}$ and it is reported in the main text in equation (16). The expressions for the average heat currents can be derived analogously and are reported in equation (IV). As a final remark we quote the expressions of the fluctuations for the separate power contributions obtained at the lowest perturbative order by following similar steps as outlined above:
$D_{P_{l}}=$ $\Omega^{2}f_{l}^{2}(\Omega)\int_{-\infty}^{+\infty}\frac{{\rm d}\omega}{4\pi m}{\cal J}_{1}(\omega+\Omega)N(\omega,\Omega)$
@@
×$\times\!\bigg{[}f_{A}^{2}(\Omega){\chi_{2}^{(A,A)}}^{\prime\prime}\!(\omega)\!+\!f_{B}^{2}(\Omega){\chi_{2}^{(B,B)}}^{\prime\prime}\!(\omega)$
$+\!2\cos(\phi)f_{A}(\Omega)f_{B}(\omega){\chi_{2}^{(A,B)}}^{\prime\prime}\!(\omega)\bigg{]}.$ (48)
@@ Appendix C Proof of $Q_{P}\geq 2$
see equation (7). We start observing that equation (16) can be rewritten as
@@
$\chi_{\mathrm{eff}}^{\prime\prime}(\omega)=$ $f_{A}^{2}(\Omega){\chi_{2}^{(A,A)}}^{\prime\prime}\!(\omega)\!+\!f_{B}^{2}(\Omega){\chi_{2}^{(B,B)}}^{\prime\prime}\!(\omega)$
$+2\cos(\phi)f_{A}(\Omega)f_{B}(\Omega){\chi_{2}^{(A,B)}}^{\prime\prime}\,,$ (51)
and we can similarly rewrite the expression of the total power fluctuations of equation (B). Finally, the entropy production rate $\dot{S}=-J_{1}/T_{1}+(P+J_{2})/T_{2}$ using equations (16)-(IV) reads
@@
By plugging equations (B), (50) and (52) into equation (49) and performing straightforward algebra, the following condition is found:
@@
and where $\lambda_{\omega}=\frac{\omega}{2T_{2}}-\frac{\omega+\Omega}{2T_{1}}$. We will now show that the condition of equation (53) can never be satisfied.
@@
Clearly, since $\lambda_{\omega}\coth(\lambda_{\omega})\geq 1$ one has $G(\omega,\omega^{\prime})\geq\left(\lambda_{\omega}-\lambda_{\omega^{\prime}}\right)^{2}\geq 0$ and therefore the sign of equation (C) is determined by the other factors in its integrand.
@@
The denominator of equation (57) is strictly positive, while the identity
@@
One then concludes that the sign of the integrand in equation (C) is given by
@@
$\mathcal{N}(\omega)=$ $f_{B}^{2}(\Omega)\left(\omega^{2}-\omega_{A}^{2}\right)+f_{A}^{2}(\Omega)\left(\omega^{2}-\omega_{B}^{2}\right)+$
$+$ $2\cos(\phi)\prod_{l=A,B}f_{l}(\Omega)\left(\omega^{2}-\omega_{l}^{2}\right)$ (62)
and where $\mathcal{D}(\omega)$ is given in equation (10). Here, $\mathcal{N}(\omega)$ is a quadratic form in $\omega^{2}$ with discriminant $-4f_{A}^{2}(\Omega)f_{B}^{2}(\Omega)(\omega_{A}^{2}-\omega_{B}^{2})^{2}\sin^{2}(\phi)\leq 0$, which proves that $\mathcal{N}(\omega)\geq 0$, whence we conclude that $\omega\chi_{\mathrm{eff}}^{\prime\prime}(\omega)\geq 0$. This finally shows that the integrand of equation (C) is non-negative and thus $I\geq 0$, contradicting the condition stated in equation (53). Thus, it is proven that $Q_{P}\geq 2$.
@@ Appendix D Asymptotic expressions
$P_{l}\sim\delta P_{l}\to\Omega f_{A}(\Omega)f_{B}(\Omega){\delta\alpha}^{({\rm ad})}_{l},$ (65)
@@
The associated fluctuations start as
$D^{({\rm ad})}_{P_{l}}$ $=\Omega^{2}f_{l}^{2}(\Omega)\int_{-\infty}^{+\infty}\frac{{\rm d}\omega}{4\pi m}{\cal J}_{1}(\omega)N(\omega,0)$
@@
$P\to\Omega{\cal J}_{1}(\Omega)\sum_{l=A,B}f_{l}^{2}(\Omega)\alpha_{l}^{({\rm dia})},$ (68)
@@
$\delta P_{l}\to\Omega f_{A}(\Omega)f_{B}(\Omega){\cal J}_{1}(\Omega)\left[\delta\alpha_{l}^{({\rm dia})}\frac{\Omega}{\omega_{c}}+\delta\beta^{({\rm dia})}_{l}\right]~{},$ (70)
@@
$\delta\alpha^{({\rm dia})}_{A/B}$ =±$=\pm\int_{-\infty}^{+\infty}\frac{{\rm d}\omega}{4\pi m}\coth\left(\frac{\omega}{2T_{2}}\right){\chi_{2}^{(A,B)}}^{\prime\prime}(\omega)~{}$
@@
$D^{({\rm dia})}_{P_{l}}=$ $\Omega^{2}f_{l}^{2}(\Omega){\cal J}_{1}(\Omega)\coth\left(\frac{\Omega}{2T_{1}}\right)$
@@
×$\times\int_{-\infty}^{+\infty}\frac{{\rm d}\omega}{4\pi m}\coth\left(\frac{\omega}{2T_{2}}\right){\chi_{2}^{(l,l)}}^{\prime\prime}(\omega).$ (72)
Appendix E TUR for the $l=B$ subsystem
Here we report results concerning the thermodynamic uncertainty relation and the associated power for the $l=B$ subsystem. These results have been obtained in the same parameter regions as in figure 3 of the main text. Also in this case regions of TUR violation are present both in the adiabatic and diabatic regime as for $l=A$ discussed in the main text. However, to these regions always correspond a positive $P_{B}>0$ and hence no useful resource can be obtained from $l=B$ in this case (see figure 7).
Figure 7: Quantifying fluctuations of $P_{B}$ via the trade-off parameter $Q_{P_{B}}$ in the isothermal regime. Density plots of $Q_{P_{B}}$ (a) and associated power $P_{B}$ (c) as a function of $\Omega/\omega_{A}$ and $T/\omega_{A}$ in the large cut-off regime $\omega_{c}=1000\omega_{A}$ and for $r=0$. Panels (b,d) report the same quantities as in panels (a,c) respectively but for a small cut-off $\omega_{c}=1.2\omega_{A}$. All results have been obtained in the strong damping regime $\gamma_{2}=100\omega_{A}$ and other parameters as in figure 3.
Appendix F Work-to-work conversion efficiency
In this appendix we consider the work-to-work conversion efficiency in the isothermal regime when $P_{A}<0$. We remind its definition:
$\eta=\frac{|P_{A}|}{P_{B}}=\frac{|P_{A}^{(0)}+f_{B}(\Omega)\delta P_{A}|}{f_{B}^{2}(\Omega)P_{B}^{(0)}-f_{B}(\Omega)\delta P_{A}}~{},$ (73)
where, we recall, $f_{B}(\Omega)=(\Omega/\omega_{c})^{r}$.
In figure 8(a) we show the work-to-work conversion efficiency in the case of $r=0$ and in the two opposite case of weak and strong damping $\gamma_{2}$. This quantity is plotted as a function of the external frequency $\Omega$, showing that in the strong damping regime (when synchronization is established) optimal efficiency close to unity is achieved in the diabatic regime with sizeable output power. In figure 8(b), instead, we consider the work-to-work conversion efficiency at strong damping $\gamma_{2}=100\omega_{A}$ but for different values of the parameter $r$. Indeed we recall that both subsystem power in general will depend on the shape of $f_{B}(\Omega)=(\Omega/\omega_{c})^{r}$. In particular, in the diabatic regime one gets the asymptotic behaviour
$\eta\to\bigg{[}1+\bigg{(}\frac{\Omega}{\omega_{c}}\bigg{)}^{r-1}\frac{\alpha_{B}^{{\rm dia}}}{|\delta\alpha_{A}^{{\rm dia}}|}\bigg{]}^{-1}$ (74)
It is clear that if $r>1$ the efficiency in the diabatic regime tends to vanish, therefore one should consider $10\leq r<$ 1 to still obtain a high efficiency.
Figure 8: Work-to-work conversion efficiency as a function of external frequency. (a) Work-to-work conversion efficiency in the weak ($\gamma_{2}=0.01\omega_{A}$) and strong ($\gamma_{2}=100\omega_{A}$) damping regime in the case $r=0$. Other parameters are $\gamma_{1}=0.001\omega_{A}$, $\omega_{c}=1.2\omega_{A}$, $T=0.1\omega_{A}$, and $\phi=\pi/2$. (b) Same quantity as in panel (a) but for different values of $r$ at strong damping ($\gamma_{2}=100\omega_{A}$).
Figure 9: Quantifying fluctuations of $P_{A}$ via the trade-off parameter $Q_{P_{A}}$ at weak damping. Density plots of $Q_{P_{A}}$ (a) and associated power $P_{A}$ (b) as a function of $\Omega/\omega_{A}$ and $T/\omega_{A}$ in the weak damping regime $\gamma_{2}=0.01\omega_{A}$ and for $r=0$. Here $\gamma_{1}=0.001\omega_{A}$, the cut-off is $\omega_{c}=1.2\omega_{A}$. Other parameters as in figure 3.
Appendix G TUR at weak damping $\gamma_{2}$
Here we present analogous results regarding the TUR quantifier $Q_{P_{A}}$ and the associated average power $P_{A}$ in the case of weak damping $\gamma_{2}\ll\omega_{l}$.
In particular, in figure 9 we set $\gamma_{2}=0.01\omega_{A}$, and we show density plots for both quantities in the $\Omega-T$ plane. As one can see, TUR violation in this case is present only in the low frequency and low temperature regime (left bottom corner of the plot in panel (a)). In this weak damping regime synchronization is absent and no TUR violation with sizeable power (in the diabatic regime) can be achieved.
Appendix H Frequency threshold for TUR violation at finite temperature gradient
In figure 10 we report the behaviour of the frequency threshold above which a TUR violation $Q_{P_{A}}<2$ is obtained in presence of a temperature gradient. Here, one can see that a minimum of $\Omega_{\rm th}$ is attained at large positive values of $\Delta T/\bar{T}$ which lowers as $\bar{T}$ increases.
Figure 10: Frequency threshold for TUR violation at finite temperature gradient. Threshold driving frequency $\Omega_{\rm th}$ such that $Q_{P_{A}}<2$ for $\Omega>\Omega_{\rm th}$ as a function of $\Delta T/\bar{T}$ in the case $r=0$ and for different values of $\bar{T}$. Other parameters as in figure 5.